OpenStudy (anonymous):

Factor each of the following. a) sin^2θ-cos^2θ b) tan^2θ-cot^2θ c) sec^2a-1

hartnn (hartnn):

\(\huge a^2-b^2=(a+b)(a-b)\)

OpenStudy (lgbasallote):

another hint... \[\cos (2 \theta) = \cos^2 \theta - \sin ^2 \theta\]

OpenStudy (lgbasallote):

but i guess that's too trig

OpenStudy (anonymous):

..

OpenStudy (anonymous):

LOL